# ScaffoldGPT: A Scaffold-based GPT Model for Drug Optimization

### v2

@@ 5.1 Experimental configuration
Baselines. We compare against baseline models: DrugImprover (Liu et al., 2023a), which utilizes an LSTM-based generator with APO fine-tuning; Molsearch (Sun et al., 2022), a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization; MIMOSA (Fu et al., 2021), a sampling-based method leveraging graph-based molecular optimization; and DrugEx v3 (Liu et al., 2023b), which leverages transformer-based reinforcement learning for scaffold-guided drug optimization. Additionally, we incorporate the model proposed by He et al. (2021; 2022), Loeffler et al. (2024), which trains a transformer to adhere to the Matched Molecular Pair (MMP) guidelines (Kenny and Sadowski, 2005, Tyrchan and Evertsson, 2017). Specifically, given a set $\{\{X,Y,Z\}\}$, where $X$ represents the source molecule, $Y$ denotes the target molecule, and $Z$ signifies the property change between $X$ and $Y$, the model learns a mapping from $\{X,Z\}\in\mathcal{X}\times\mathcal{Z}\implies Y\in\mathcal{Y}$ during training. Here, $\mathcal{X}\times\mathcal{Z}$ denotes the input space, and $\mathcal{Y}$ denotes the target space. They defined six different types of property changes for $Z$, including MMP for user-specified alterations, various similarity thresholds, and scaffold-based modifications where molecules share the same scaffold or a generic scaffold. More specifically,
@@
Scaffold (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.609 $\pm$ 0.001 0.718 $\pm$ $0$8.508 $\pm$ 0.026 0.711 $\pm$ 0.000 2.627 $\pm$ 0.002 2.803 $\pm$ 0.010 0.735 $\pm$ 0.002
Scaffold Generic (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.625 $\pm$ 0.001 0.722 $\pm$ $0$8.544 $\pm$ 0.009 0.722 $\pm$ 0.002 2.551 $\pm$ 0.010 2.898 $\pm$ 0.005 0.768 $\pm$ 0.004
@@
Docking ()($\downarrow$)( ↓ ) -10.031 -9.258 -11.478 $\checkmark$ -11.474 $\checkmark$ -11.087 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.646 0.624 0.762 $\checkmark$ 0.774 $\checkmark$ 0.762 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 2.390 2.417 2.298 $\checkmark$ 2.257 $\checkmark$ 2.356 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 2.590 2.680 $\checkmark$ 4.007 $\checkmark$ 2.893 $\checkmark$ 4.007 $\checkmark$
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.618 0.589 0.759 $\checkmark$ 0.753 $\checkmark$ 0.754 $\checkmark$
@@ 6 Conclusion
We have introduced ScaffoldGPT, a novel framework for drug optimization. This framework incorporates a unique scaffold-based GPT design, a three-step optimization process, a two-phase incremental training method, and a novel Top-N decoding strategy that facilitates controlled reward-guided generation using GPT. To showcase the superior performance of ScaffoldGPT, we conduct evaluations on real-world viral and cancer-related datasets to compare it against eight competing baselines. Our results demonstrate that ScaffoldGPT surpasses all baselines across the majority of performance metrics, underscoring its efficacy. Our work highlights ScaffoldGPT’s effectiveness in drug optimization, as evidenced by enhancements in various pharmaceutical properties. Currently, ScaffoldGPT is limited to handling molecules in SMILES format. We are working to expand ScaffoldGPT’s capabilities to accommodate a broader range of drug representation formats as a future direction.
@@ A.3 BPE Tokenization
BPE constructs a robust vocabulary by iteratively merging the most frequent token pairs, effectively capturing common subword units for more efficient and flexible text representation. The resulting vocabulary comprises 3,152 tokens and includes special tokens as well. For instance, the sequence <L> is tokenized into three separate tokens: $<$, L, and $>$. The tokenizer was trained on 10 million molecules from the ZINC dataset, ensuring comprehensive coverage of chemical elements.
@@ A.5 Drug Optimization illustration on COVID benchmark
Docking ()($\downarrow$)( ↓ ) -9.748 -10.184 $\checkmark$ -10.187 $\checkmark$
Druglikeness ()($\uparrow$)( ↑ ) 0.839 0.840 $\checkmark$ 0.847 $\checkmark$
Synthesizability ()($\downarrow$)( ↓ ) 5.631 1.983 $\checkmark$ 2.199 $\checkmark$
Solubility ()($\uparrow$)( ↑ ) 0.192 5.079 $\checkmark$ 3.906 $\checkmark$
@@
Avg Norm Reward ()($\uparrow$)( ↑ ) 0.398 0.688 $\checkmark$ 0.694 $\checkmark$
@@ A.7 Hyperparameters and architectures
Table 6 and 7 provides a list of hyperparameter settings we used for our experiments.
For APO finetuning and experimentation, 1280 molecules were selected from each of the RTCB and 3CLPro datasets, with docking scores ranging from -14 to -6. This range is based on (Liu et al., 2024b).
@@
Moreover, when the generated SMILES is invalid, indicating that the reward $R_{c}$ cannot be calculated, we have two options: either directly subtract the reward of the original SMILES (i.e., $-R_{c}(X)$), or consider the advantage preference as zero instead.

### v3

@@ 5.1 Experimental configuration
Baselines. We compare against baseline models: DrugImprover (Liu et al., 2023a), which utilizes an LSTM-based generator with APO fine-tuning; Molsearch (Sun et al., 2022), a Monte Carlo tree search (MCTS)-driven approach for molecular generation and optimization; MIMOSA (Fu et al., 2021), a sampling-based method leveraging graph-based molecular optimization; and DrugEx v3 (Liu et al., 2023d), which leverages transformer-based reinforcement learning for scaffold-guided drug optimization. Additionally, we incorporate the model proposed by He et al. (2021, 2022); Loeffler et al. (2024), which trains a transformer to adhere to the Matched Molecular Pair (MMP) guidelines (Kenny and Sadowski, 2005; Tyrchan and Evertsson, 2017). Specifically, given a set $\{\{X,Y,Z\}\}$, where $X$ represents the source molecule, $Y$ denotes the target molecule, and $Z$ signifies the property change between $X$ and $Y$, the model learns a mapping from $\{X,Z\}\in\mathcal{X}\times\mathcal{Z}\implies Y\in\mathcal{Y}$ during training. Here, $\mathcal{X}\times\mathcal{Z}$ denotes the input space, and $\mathcal{Y}$ denotes the target space. They defined six different types of property changes for $Z$, including MMP for user-specified alterations, various similarity thresholds, and scaffold-based modifications where molecules share the same scaffold or a generic scaffold. More specifically,
@@
Scaffold (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.609 $\pm$ 0.001 0.718 $\pm$ 0.000 -8.508 $\pm$ 0.026 0.711 $\pm$ 0.000 2.627 $\pm$ 0.002 2.803 $\pm$ 0.010 0.735 $\pm$ 0.002
Scaffold Generic (Loeffler et al., 2024) 0.998 $\pm$ 0.001 0.625 $\pm$ 0.001 0.722 $\pm$ 0.000 -8.544 $\pm$ 0.009 0.722 $\pm$ 0.002 2.551 $\pm$ 0.010 2.898 $\pm$ 0.005 0.768 $\pm$ 0.004
@@
Docking $(\downarrow)$ -10.031 -9.258 -11.478 $\checkmark$ -11.474 $\checkmark$ -11.087 $\checkmark$
Druglikeness $(\uparrow)$ 0.646 0.624 0.762 $\checkmark$ 0.774 $\checkmark$ 0.762 $\checkmark$
Synthesizability $(\downarrow)$ 2.390 2.417 2.298 $\checkmark$ 2.257 $\checkmark$ 2.356 $\checkmark$
Solubility $(\uparrow)$ 2.590 2.680 $\checkmark$ 4.007 $\checkmark$ 2.893 $\checkmark$ 4.007 $\checkmark$
Avg Norm Reward $(\uparrow)$ 0.618 0.589 0.759 $\checkmark$ 0.753 $\checkmark$ 0.754 $\checkmark$
@@
Top-N selects a higher‑reward alternative missed by Top-K.
These two SMILES strings are nearly identical, but TOP-N selects a slightly different generation path—preferring ‘c’ and ‘c1C’ over ‘c(C)c1’—which results in a higher average reward. This demonstrates how TOP-N can identify token-level alternatives that improve the optimization objective while preserving scaffold structure.
Method Generated SMILES
Avg.
Normalized Reward
Top-K Cc1ccc(NC(=O)COC(=O)C2C[C@H]3CCCC@HC3=O)c(C)c1 0.579
Top-N Cc1cccc(NC(=O)COC(=O)C2C[C@H]3CCCC@HC3=O)c1C 0.614
Table 4: Comparison of generated SMILES and average normalized rewards across methods.
Ablation study of Top-N vs Top-P, Top-K and Top-PK strategies.
@@ 6 Conclusion
We have introduced ScaffoldGPT, a novel framework for drug optimization. This framework incorporates a unique scaffold-based GPT design, a three-stage optimization process, a two-phase incremental pre-training approach, and a novel Top-N decoding strategy that facilitates controlled reward-guided generation using GPT. To showcase the superior performance of ScaffoldGPT, we conduct evaluations on real-world viral and cancer-related datasets to compare it against eight competing baselines. Our results demonstrate that ScaffoldGPT surpasses all baselines across the majority of performance metrics, underscoring its efficacy. Our work highlights ScaffoldGPT’s effectiveness in drug optimization, as evidenced by enhancements in various pharmaceutical properties. Currently, ScaffoldGPT is limited to handling molecules in SMILES format. We are working to expand ScaffoldGPT’s capabilities to accommodate a broader range of drug representation formats as a future direction.
Acknowledgements
This work is supported in part by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.
@@ A.3 BPE Tokenization
BPE constructs a robust vocabulary by iteratively merging the most frequent token pairs, effectively capturing common subword units for more efficient and flexible text representation. The resulting vocabulary comprises 3,152 tokens and includes special tokens as well. For instance, the sequence ¡L¿ is tokenized into three separate tokens: <<, L, and >>. The tokenizer was trained on 10 million molecules from the ZINC dataset, ensuring comprehensive coverage of chemical elements.
@@ A.5 Drug Optimization illustration on COVID benchmark
Docking $(\downarrow)$ -9.748 -10.184 $\checkmark$ -10.187 $\checkmark$
Druglikeness $(\uparrow)$ 0.839 0.840 $\checkmark$ 0.847 $\checkmark$
Synthesizability $(\downarrow)$ 5.631 1.983 $\checkmark$ 2.199 $\checkmark$
Solubility $(\uparrow)$ 0.192 5.079 $\checkmark$ 3.906 $\checkmark$
@@
Avg Norm Reward $(\uparrow)$ 0.398 0.688 $\checkmark$ 0.694 $\checkmark$
@@ A.7 Hyperparameters and architectures
Table 7 and 8 provides a list of hyperparameter settings we used for our experiments.
For APO finetuning and experimentation, we have two different set of 1280 molecules that were selected from each of the RTCB and 3CLPro datasets, with docking scores ranging from -14 to -6. This range is based on (Liu et al., 2024c).
@@
Moreover, when the generated SMILES is invalid, indicating that the reward $R_{c}$ cannot be calculated, we have two options: either directly subtract the reward of the original SMILES (i.e., $R_{c}(X)$), or consider the advantage preference as zero instead.